# Improbable Bigrams Expose Vulnerabilities of Incomplete Tokens in Byte-Level Tokenizers

### v1

@@ Abstract
Tokenization is a crucial step that bridges human-readable text with model-readable discrete tokens. However, recent studies have revealed that tokenizers can be exploited to elicit unwanted model behaviors. In this work, we investigate incomplete tokens, i.e., undecodable tokens with stray bytes resulting from byte-level byte-pair encoding (BPE) tokenization. We hypothesize that such tokens are heavily reliant on their adjacent tokens and are fragile when paired with unfamiliar tokens. To demonstrate this vulnerability, we introduce improbable bigrams: out-of-distribution combinations of incomplete tokens designed to exploit their dependency. Our experiments show that improbable bigrams are significantly prone to hallucinatory behaviors. Surprisingly, alternative tokenizations of the same phrases result in drastically lower rates of hallucination (93% reduction in Llama3.1). We caution against the potential vulnerabilities introduced by byte-level BPE tokenizers, which may impede the development of trustworthy language models.
@@
Eugene Jang1     Kimin Lee2     Jin-Woo Chung1     Keuntae Park1     Seungwon Shin2 1S2W Inc.    2KAIST 1{genesith,jwchung,keuntae.park}@s2w.inc    2{kiminlee,claude}@kaist.ac.kr
@@ 2 BPE Tokenization
Tokenizer Vocab Size Incomplete Tokens Incomplete Bigrams
@@ 3 Incomplete Tokens and Improbable Bigrams
{CJK*}
UTF8gbsn
Byte-level BPE tokenizers are agnostic not only to morphemes but also to character boundaries. Training with the BPE algorithm on texts where characters are encoded with multiple bytes often results in the formation of incomplete tokens. Incomplete tokens contain stray bytes (like <0x9F> in the token ‘<0x9F>能’) that do not correspond to recognizable characters. These incomplete tokens constitute a small but significant portion of the byte-level vocabulary. Table 1 presents the number of discovered incomplete tokens in the vocabularies of various models employing byte-level tokenizers (see Appendix A for details).
@@
Table 2: Prompt templates used to test phrase-level hallucinations of target phrases. At test time, {Phrase} is replaced with the tested phrase. If the model fails to repeat the phrase in all three prompts, we consider the phrase to be hallucinatory.
@@
While incomplete tokens are indistinguishable from complete tokens from the model’s perspective during training, their structure can influence how they are trained in subtle ways. First, incomplete tokens can only occur alongside other incomplete or single-byte tokens, as the “stray bytes” of incomplete tokens require additional bytes to form a viable character. This restriction limits the range of tokens they can be associated with. Second, the potential characters that these stray bytes can form may have little to no semantic overlap, leading to a representation that is inherently ambiguous and heavily reliant on neighboring tokens for disambiguation. We hypothesize that these factors exacerbate one another, causing incomplete tokens to become overly dependent on their context and brittle in response to unfamiliar adjacent tokens.
To demonstrate this vulnerability, we analyze the structure of incomplete tokens, categorizing those that end with stray bytes as prefixes and those that begin with stray bytes as suffixes. We further label prefixes and suffixes by the number of stray bytes they require to form a complete character, to find potential bigram combinations. We create adversarial combinations of these tokens, termed improbable bigrams, as illustrated in Figure 2. These bigrams exploit the dependency of two incomplete tokens to complete each other. We note that not all combinations are viable due to tokenization rules and merge priorities. Table 1 lists the number of all possible legal bigrams created from incomplete tokens across different tokenizers. To identify bigrams that are particularly out-of-distribution, we use the heuristic of multilinguality: if a bigram combines to form scripts from different languages, we assume it is highly unlikely to have been encountered during training. In our experiments, we demonstrate that improbable bigrams are prone to hallucinations.
@@ 4 Hallucination Experiments
Models Improbable Bigrams Baseline Bigrams Llama 3.1 43/100 (43%) 0/100 (0%) Exaone 79/100 (79%) 26/100 (26%) Qwen2.5 38/100 (38%) 2/100 (2%) Mistral-Nemo 52/71 (73%) 0/71 (0%) Command-R 56/100 (56%) 9/100 (9%)
@@
Evaluation Prompts We assess each model’s ability to accurately process prompts containing improbable bigram phrases. We use three specific prompt templates designed to induce the model to repeat a target phrase from the input, regardless of the phrase’s nonsensical or unusual nature, as detailed in Table 2. We consider a phrase to have induced hallucinations only if the model fails to correctly repeat it across all three prompt templates.
Improbable Bigrams For each model, we construct improbable bigrams using the procedure outlined in Section 3. Noting that undertrained tokens often lead to hallucinations, we focus our experiments on well-trained incomplete tokens to isolate other factors. Employing the embedding heuristic methodology from Land and Bartolo (2024), we rank all tokens by their degree of training and select only the incomplete tokens in the upper half of this ranking. We then create up to 100 improbable bigrams for each model for our experiments.
Baselines To better demonstrate the fragility of improbable bigrams, we create a baseline of bigrams constructed with complete tokens. For each improbable bigram, we create a complete token counterpart by replacing the prefix and suffix with similarly-trained tokens. That is, from an initial incomplete token that is the i-th token in the ranked vocabulary, we find a complete token that is slightly less trained (e.g. the (i-1)-th token). This method ensures that each improbable bigram is compared to a baseline of complete bigrams with a similar level of undertraining.
@@
5 Dehallucinations via Alternative Tokenization
Models Incomplete Token Hallucinations Alternative Tokenization Hallucinations Llama 3.1 0.43 0.03 (↓93%) Exaone 0.79 0.52 (↓34%) Qwen2.5 0.38 0.14 (↓63%) Mistral-Nemo 0.73 0.01 (↓98%) Command-R 0.56 0.57
@@ 5 Dehallucinations via Alternative Tokenization
While we have hypothesized that improbable bigram hallucinations are caused by incomplete tokens, another possible explanation for their hallucinations is that they are simply just out-of-distribution. To test this hypothesis, we repeat the experiments using the same phrases, but with a different tokenization method that avoids incomplete tokens. Instead of tokenizing the bigram phrase according to the tokenizer, we avoid character boundary-crossing tokenization using pre-segmentation. {CJK*}UTF8ipxm For instance, the improbable bigram phrase “サーミ能” is normally tokenized into incomplete tokens ‘サー<0xE3><0x83>’ and ‘<0x9F>能’. However, it can be presegmented to isolate the character formed from stray bytes (‘ミ’). We split the string into three parts (‘サー’, ‘ミ’, and ‘能’) and tokenize each separately. Afterwards, the tokens can be appended together to create an alternative token representation of the same phrase.
Compared to improbable bigrams, alternative tokenization sequences can be suboptimal in two important ways. First, these sequences cannot be naturally generated by the tokenizer, ensuring that they are out-of-distribution for the model. Second, the sequences consist of more than two tokens, requiring the model to recall more tokens correctly to repeat the phrase accurately. If the out-of-distribution hypothesis holds, we would expect these phrases to induce more hallucinations with this alternative tokenization.
Table 4 shows results of the experiment repeated using the alternative tokenization scheme. Surprisingly, we find that these token sequences, which are both unnatural and longer, hallucinate significantly less than the original tokenization for most models. These results support our hypothesis that the hallucinatory behaviors are being caused by the incomplete tokens.
@@ 6 Conclusion
Our findings on improbable bigrams causing hallucinations were consistent in all 5 models. However, our secondary experiment on alternate tokenization did not cause improvements for the Cohere Command-R model. We hypothesize that the training of the tokens may have been different due to the model’s significantly larger vocabulary size, which is approximately double the other studied models. However, a more detailed investigation to verify this is left to future work. There is a possibility that these findings may be influenced by other unknown dependencies relating to tokenizer parameters and training parameters.
@@ Appendix A Incomplete Bigram Construction
UTF-8 characters have a flexible number of bytes. Each multi-byte character is comprised of a starting byte and continuation bytes. Starting bytes indicate how many bytes the character is composed of. For example, the starting byte of a 3-byte character implies the following two characters will be continuation bytes. A more detailed explanation of the UTF-8 protocol and byte structures can be found in Limisiewicz et al. (2024).

### v2

@@ Abstract
Tokenization is a crucial step that bridges human-readable text with model-readable discrete tokens. However, recent studies have revealed that tokenizers can be exploited to elicit unwanted model behaviors. In this work, we investigate incomplete tokens, i.e., undecodable tokens with stray bytes resulting from byte-level byte-pair encoding (BPE) tokenization. We hypothesize that such tokens are heavily reliant on their adjacent tokens and are fragile when paired with unfamiliar tokens. To demonstrate this vulnerability, we introduce improbable bigrams: out-of-distribution combinations of incomplete tokens designed to exploit their dependency. Our experiments show that improbable bigrams are significantly prone to hallucinatory behaviors. Surprisingly, the same phrases have drastically lower rates of hallucination (90% reduction in Llama3.1) when an alternative tokenization is used. We caution against the potential vulnerabilities introduced by byte-level BPE tokenizers, which may introduce blind spots to language models.
@@
Eugene Jang1thanks:  Work performed while at S2W Inc.    Kimin Lee2    Jin-Woo Chung3    Keuntae Park3    Seungwon Shin2 1Northeastern University  2KAIST  3S2W Inc. 1jang.eu@northeastern.edu2{kiminlee,claude}@kaist.ac.kr 3{jwchung,keuntae.park}@s2w.inc
@@ 2 BPE Tokenization
Tokenizer
Vocab
Size
Incomplete
Tokens
Incomplete
Bigrams
@@ 3 Incomplete Tokens and Improbable Bigrams
Byte-level BPE tokenizers are agnostic not only to morphemes but also to character boundaries. Training with the BPE algorithm on texts where characters are encoded with multiple bytes (e.g., Cyrillic (2 bytes), CJK Unified Ideographs (3 bytes)) often results in the formation of incomplete tokens. Incomplete tokens contain stray bytes (like <0x9F> in the token ‘<0x9F>能’) that do not correspond to recognizable characters. These incomplete tokens constitute a small but significant portion of the vocabulary in byte-level models. Table 1 presents the number of discovered incomplete tokens in the vocabularies of various models employing byte-level tokenizers (see Appendix A for details).
@@
Repeat Repeat this phrase exactly: ‘‘{Phrase}’’
@@
Table 2: Prompt templates used to test phrase-level hallucinations of target phrases. At test time, {Phrase} is replaced with the tested phrase. If the model fails to repeat the phrase in all four prompts, we consider the phrase to be hallucinatory.
@@
While incomplete tokens are indistinguishable from complete tokens from the model’s perspective during training, their structure can influence how they are trained in subtle ways. First, incomplete tokens can only occur alongside other incomplete or single-byte tokens, as the “stray bytes” of incomplete tokens require additional bytes to form a viable character. This restriction limits the range of tokens they can be associated with. Second, the potential characters that stray bytes can form into may have little to no semantic overlap, leading to a representation that is inherently ambiguous and heavily reliant on neighboring tokens for disambiguation. We hypothesize that these factors exacerbate one another, causing incomplete tokens to become overly dependent on their context and brittle in response to unfamiliar adjacent tokens.
To demonstrate this vulnerability, we create phrases that exploit incomplete tokens to cause hallucinations. First, we analyze the structure of incomplete tokens, categorizing those that end with stray bytes as prefixes and those that begin with stray bytes as suffixes. We further label prefixes and suffixes by the number of stray bytes they require to form a complete character, to find potential bigram combinations. We create adversarial combinations of these tokens, termed improbable bigrams, as illustrated in Figure 2. These bigrams exploit the dependency of two incomplete tokens to complete each other. We note that not all combinations are viable due to tokenization rules and merge priorities. Table 1 lists the number of all possible legal bigrams created from incomplete tokens across different tokenizers. To identify bigrams that are particularly out-of-distribution, we use the heuristic of multilinguality. When a bigram forms a string, we check the Unicode script type of each character. If the string contains characters from different Unicode scripts, we assume it is highly unlikely to have been encountered during training. In our experiments, we demonstrate that improbable bigrams are prone to hallucinations.
@@ 4 Hallucination Experiments
Models Improbable Bigrams Baseline Bigrams Llama 3.1 48/100 (48%) 0/100 (0%) Exaone 77/100 (77%) 20/100 (20%) Qwen2.5 33/100 (33%) 0/100 (0%) Mistral-Nemo 52/71 (73%) 1/71 (1%) Command-R 49/100 (49%) 8/100 (8%)
@@
Evaluation Prompts We assess each model’s ability to accurately process prompts containing improbable bigram phrases. We use four prompt templates designed to induce the model to repeat a target phrase from the input, regardless of the phrase’s nonsensical or unusual nature, as detailed in Table 2. Our goal is to find phrases that models cannot process for any prompt. Therefore, we consider a phrase to have induced hallucinations only if the model fails to correctly repeat it across all four prompt templates.
Incomplete Token Selection For each model, we construct improbable bigrams using the procedure outlined in Section 3. Noting that undertrained tokens often lead to hallucinations, we focus our experiments on well-trained tokens to isolate confounding factors. To do this, we employ the embedding heuristics in Land and Bartolo (2024) to detect whether a token in the vocabulary is undertrained. Specifically, the heuristics estimate the degree of training for each token by taking L2 norms and cosine distances in the embedding matrix. We use these metrics to order the entire vocabulary, and discard the lower (undertrained) half from our experiments. From the upper half (well-trained half) of the vocabulary, we identify the incomplete tokens and create up to 100 improbable bigrams for each model. Analysis of the languages used in the improbable bigrams (Appendix B) show language-pair distributions differ widely between models.
Baselines To better demonstrate the fragility of improbable bigrams, we create a baseline of bigrams constructed with complete tokens. For each improbable bigram, we create a complete token counterpart by replacing the prefix and suffix with similarly-trained tokens. That is, if the initial incomplete token is the i-th token in the ranked vocabulary, we find a complete token that is slightly less trained (e.g. the (i-1)-th token). This method ensures that each improbable bigram is compared to a baseline of complete bigrams with a similar level of undertraining.
@@
5 Alternative Tokenization
Models Incomplete Token Hallucinations Alternative Tokenization Hallucinations Llama 3.1 0.48 0.05 (↓90%) Exaone 0.77 0.50 (↓35%) Qwen2.5 0.33 0.12 (↓64%) Mistral-Nemo 0.73 0.01 (↓98%) Command-R 0.49 0.55
@@ 5 Alternative Tokenization
Our previous experiments demonstrate that improbable bigram phrases are difficult for models to process. We conduct further experiments to attribute this difficulty to the incomplete tokens. We repeat the experiments, but pre-segment the target phrase to avoid character boundary-crossing tokenization. For instance, the improbable bigram phrase “サーミ能” is normally tokenized into incomplete tokens “サー<0xE3><0x83>” and “<0x9F>能”. However, it can be presegmented to isolate the character formed from stray bytes (“ミ”). We split the string into three parts (“サー”, “ミ”, and “能”) and tokenize each separately. Afterwards, the tokens can be appended together to create an alternative token representation of the same phrase.
Compared to improbable bigrams, the alternative tokenization sequences can be suboptimal in two important ways. First, these sequences cannot be generated by the tokenizer, ensuring that they are out-of-distribution for the model. Second, the sequences consist of more than two tokens, requiring the model to recall more tokens correctly to repeat the phrase accurately.
Table 4 shows results of the experiment repeated using the alternative tokenization scheme. Despite the suboptimalities of alternative tokenization, most models, with the exception of Command-R (discussed in Appendix B) show significantly higher performance when alternative tokenization is used. The same phrases tend to hallucinate significantly less when alternative tokenization is used, suggesting that the hallucinatory behaviors are being caused by the incomplete tokens.
6 Discussion
Risks of Unrepeatability We have used a set of repetition-inducing prompts as a simple task that identifies phrases that cause model failures. The repetition failures, stemming from incomplete tokens, are possibly indicative of other robustness failures linked to incomplete tokens. However, in certain scenarios, non-repeatability itself can cause real world harms. We highlight few such scenarios.
  • Reliable Code and Data Handling: When models interact with code or databases, they must correctly preserve variable names or fixed values. If models are unable to preserve certain phrases, they can compromise the integrity of data.
  • Adversarial unrepeatability: Adversaries may exploit unrepeatable phrases to avoid intervention from LLM agents. For instance, an attacker can set their username to an unrepeatable phrases to evade a moderator agent.
  • Model Fingerprinting: Since improbable bigrams are model-specific due to vocabulary design, they could be used to fingerprint the architecture behind a closed or anonymized LLM service. This can facilitate sophisticated targeted attacks on private LLMs (e.g. adversarial triggers optimized for the target architecture). This can also create implications for model evaluations that rely on model anonymity (e.g. Chatbot Arena).
Circumvention Strategies Our findings suggest that model developers should consider the consequences of vulnerable tokenizer design. The vulnerability of incomplete tokens can be circumvented by preventing incomplete tokens from entering the final tokenizer vocabulary (before model training begins). One method could involve vocabulary pruning Bauwens and Delobelle (2024) after tokenizer training to remove incomplete tokens. Another method would be to constrain BPE merges to respect character boundaries during tokenizer training Land and Arnett (2025). If full Unicode coverage is not deemed necessary for a certain model, developers may opt to choose tokenization at the character-level before tokenizer-training. We hope that demonstrating potential vulnerabilities of the default byte-level BPE can motivate more careful design of tokenizers.
@@ 7 Conclusion
Our findings on improbable bigrams causing hallucinations were consistent in all 5 models. However, our secondary experiment on alternate tokenization did not cause improvements for the Cohere Command-R model. It is difficult to isolate which design choices of the model contributes to this effect. One hypothesis is that the model’s significantly larger vocabulary size, which is approximately double the other studied models, may have impacted token training. We have also conducted an analysis on the languages of the incomplete bigrams as a possible explanation, but found it inconclusive. A more detailed investigation is left to future work. There is a possibility that these findings may be influenced by other unknown and undocumented dependencies relating to tokenizer parameters, training parameters, and training data.
@@ Appendix A Incomplete Bigram Construction
UTF-8 characters have a flexible number of bytes. Each multi-byte character comprises of a starting byte and continuation bytes. Starting bytes indicate how many bytes the character is composed of. For example, the starting byte of a 3-byte character implies the following two characters will be continuation bytes. A more detailed explanation of the UTF-8 protocol and byte structures can be found in Limisiewicz et al. (2024).
@@
Appendix B Role of Languages in Improbable Bigrams
We analyze the scripts of the improbable bigrams used in the experiments. Note that the language distributions of the tested improbable bigrams are not indicative of all possible improbable bigrams in each model’s vocabulary, since the bigrams from the experiments were selected from the top half of the vocabulary when ordering by token trainedness (Section 4). Figure 3 shows the distribution of languages of the improbable bigrams for each model. We observe that improbable bigrams can occur in a wide variety of languages, with higher resource multi-byte scripts such as Chinese, Korean, and Russian being the most frequent. The diversity of languages of hallucinations suggest there is no significant influence caused by languages.
Models had different distributions of language-pairs, which is influenced by both tokenizer training and model training. For instance, Exaone had 17 unique language-pairs while Command-R had only 3 unique language-pairs. Command-R’s much larger tokenizer vocabulary may have impacted the incomplete token selection. All of the 100 bigrams for Command-R contained Hebrew characters. For comparison, there are no tokens in the top-half of Llama3.1’s vocabulary that resolve to Hebrew characters. Command-R is also the only model analyzed with bigrams of the Hebrew/Korean language-pair. Bigrams of this language-pair constitute a majority of the tested improbable bigrams for Command-R, and performed worse when using alternative tokenization.
Alternative Tokenization with Hebrew/Korean The Command-R model’s improbable bigrams did not show improvements when using alternative tokenization. Figure 3 shows that alternative tokenization harmed performance for the 81 Hebrew/Korean bigrams. We investigate if the Hebrew/Korean language-pair is particularly prone to hallucinations when using alternative tokenization. We collect a total of 38 possible bigrams of this language combination from the Llama3.1 vocabulary. Repeating the experiments with Llama3.1’s 38 Hebrew/Korean bigram pairs, using alternative tokenization reduced hallucinations (20/38) compared to regular tokenization (32/38). These results suggest that Command-R’s unusual result with alternative tokenization was not caused by the language-pairing of its improbable bigrams.
(a) Improbable Bigrams
(b) Improbable Bigrams (alternative tokenization)
Figure 3: Language-pair distributions of the improbable bigrams used in experiments. Darkened colors indicate improbable bigrams that cause hallucinations.